# Statistical Properties of a Polarization Vector’s Ellipticity Angle

### v1

@@ Appendix A Confidence Limits of the Ellipticity Angle
The pdf of the EA is symmetric when the amplitude of the polarization vector is constant and the intrinsic EA is $\chi_{o}=0$. The symmetry of the pdf allows a $68\%$ confidence limit, $\sigma_{c}$, to be calculated from Equation 3 following the procedure used by Wardle & Kronberg (1974), Naghizadeh-Khouei & Clarke (1993), and Everett & Weisberg (2001) to determine the confidence limits of the PA:
@@
A different approach to calculating the confidence limits must be adopted for other values of $\chi_{o}$, because $f_{\chi}(\chi)$ is generally asymmetric in those cases. Przysucha (2024) suggested that the confidence interval of an asymmetric pdf, $f(x)$, can be calculated from its semivariances. His method is applicable to any continuous random variable with a mean, $\mu$, and standard deviation, $\sigma$. The upper semivariance, $\sigma_{+}^{2}$, is the variance of $f(x)$ calculated from values of $x$ that exceed the mean
$\sigma_{+}^{2}=\int_{\mu}^{\infty}(x-\mu)^{2}f(x)dx,$ (A3)
and the lower semivariance, $\sigma_{-}^{2}$, is the variance of $f(x)$ calculated from values of $x$ that fall below the mean
@@
Increasing the parameter $k$ in Equation A5 increases the confidence interval and the confidence limits. The confidence limits of the EA can be determined by selecting a desired value of the confidence interval, in this case $\gamma=0.6827$, and finding the value of $k$ that is the solution to
@@
The lower limit of integration in Equation A6 is $\tau_{-}=\mu_{\chi}-k\sigma_{-}$ and the upper limit is $\tau_{+}=\mu_{\chi}+k\sigma_{+}$. Once $k$ is known, the confidence limits can be calculated from the definitions of the semivariances. The value of $k$ varies with different values of $s$ and $\chi_{o}$ in $f_{\chi}(\chi)$. It is equal to $\sqrt{2}$ when the pdf is Gaussian, because the semivariances are equal to one-half the pdf’s variance and the $68\%$ confidence limit of a Gaussian distribution is equal to its standard deviation.
The resulting upper and lower confidence limits in the case of $\chi_{0}=\pi/4$ are shown by the red and blue lines, respectively, in the bottom panel of Figure 12. The dashed black line denotes the mean EA as calculated from the first moment of the pdf given by Equation 5. The filled circles that accompany the dashed black line denote the approximation to the mean EA given by Equation 7. The approximation is very good. The figure shows the confidence limits generally are not symmetric about the mean. A fit to the confidence limits calculated from Equation A6 yields an upper bound on $\chi$ that varies with $s$ as
@@
Table 1: Mean and 68% Confidence Limits of the Ellipticity Angle
@@
Example confidence limits calculated for other values of $\chi_{o}$ are tabulated in Table 1. The left column of the table lists the intrinsic EA ranging from $0^{\circ}$ to $45^{\circ}$ in increments of $\Delta\chi_{o}=5^{\circ}$. The top row of the table lists values of SNR in total polarization ranging from $s=3$ to $s=18$ in increments of $\Delta s=3$. The table entries include the mean of the pdf defined by the specified values of $s$ and $\chi_{o}$ along with the upper and lower errors determined from Equation A6. The magnitudes of the upper and lower errors for a given table entry are the same for most entries, especially when the SNR is large and the intrinsic EA is small. This is an indication that their respective pdfs are symmetric and their values of $s$ and $\chi_{o}$ are consistent with the condition $s>3/\cos(2\chi_{o})$. The magnitude of the errors for individual entries in the second column and last two rows of the table generally are not equal, indicating their respective pdfs are not symmetric.
Table 1 can serve as a look-up table for placing upper and lower limits on an EA measurement. Assuming that the measured EA is the mean EA, the table entries and the measured value of $s$ provide the errors associated with the measurement. For example, if the measured SNR is $s=3$ and the measured EA is $\chi_{m}=26.7^{\circ}$, the upper error on the measurement is $7.9^{\circ}$, and the lower error is $9.1^{\circ}$. Alternatively, analytical expressions for the upper and lower bounds as functions of $s$ for a given $\chi_{o}$ can be derived by fitting the data in Table 1 to equations similar to Equations A7 and A8. Or, a more extensive look-up table with finer resolution in $s$ and $\chi_{o}$ can be generated by numerically calculating the mean of the relevant pdf and its upper and lower bounds from Equation A6.

### v2

@@ Appendix A Confidence Limits of the EA
The pdf of the EA is symmetric when the amplitude of the polarization vector is constant and the intrinsic EA is $\chi_{o}=0$. The symmetry of the pdf allows a $68\%$ confidence limit, $\sigma_{c}$, to be calculated from Equation 3, following the procedure used by Wardle & Kronberg (1974), Naghizadeh-Khouei & Clarke (1993), and Everett & Weisberg (2001) to determine the confidence limits of the PA:
@@
A different approach to calculating the confidence limits must be adopted for other values of $\chi_{o}$, because $f_{\chi}(\chi)$ is generally asymmetric in those cases. Przysucha (2024) suggested that the confidence interval of an asymmetric pdf, $f(x)$, can be calculated from its semivariances. His method is applicable to any continuous random variable with a mean, $\mu$, and standard deviation, $\sigma$. The upper semivariance, $\sigma_{+}^{2}$, is the variance of $f(x)$ calculated from values of $x$ that exceed the mean:
$\sigma_{+}^{2}=\int_{\mu}^{\infty}(x-\mu)^{2}f(x)dx;$ (A3)
and the lower semivariance, $\sigma_{-}^{2}$, is the variance of $f(x)$ calculated from values of $x$ that fall below the mean:
@@
Increasing the parameter $k$ in Equation A5 increases the confidence interval and the confidence limits. The confidence limits of the EA can be determined by selecting a desired value of the confidence interval - in this case $\gamma=0.6827$ - and finding the value of $k$ that is the solution to
@@
The lower limit of integration in Equation A6 is $\tau_{-}=\mu_{\chi}-k\sigma_{-}$ and the upper limit is $\tau_{+}=\mu_{\chi}+k\sigma_{+}$. Once $k$ is known, the confidence limits can be calculated from the definitions of the semivariances. The value of $k$ varies with different values of $s$ and $\chi_{o}$ in $f_{\chi}(\chi)$. It is equal to $\sqrt{2}$ when the pdf is Gaussian, because the semivariances are equal to one-half the pdf’s variance, and the $68\%$ confidence limit of a Gaussian distribution is equal to its standard deviation.
The resulting upper and lower confidence limits in the case of $\chi_{0}=\pi/4$ are shown by the red and blue lines, respectively, in the bottom panel of Figure 12. The dashed black line denotes the mean EA, as calculated from the first moment of the pdf given by Equation 5. The filled circles that accompany the dashed black line denote the approximation to the mean EA given by Equation 7. The approximation is very good. The figure shows the confidence limits generally are not symmetric about the mean. A fit to the confidence limits calculated from Equation A6 yields an upper bound on $\chi$ that varies with $s$ as
@@
Table 1: Mean and 68% Confidence Limits of the EA
@@
Example confidence limits calculated for other values of $\chi_{o}$ are tabulated in Table 1. The left column of the table lists the intrinsic EA, ranging from $0^{\circ}$ to $45^{\circ}$ in increments of $\Delta\chi_{o}=5^{\circ}$. The top row of the table lists values of the SNR in total polarization, ranging from $s=3$ to $s=18$ in increments of $\Delta s=3$. The table entries include the means of the pdfs defined by the specified values of $s$ and $\chi_{o}$, along with the upper and lower errors determined from Equation A6. The magnitudes of the upper and lower errors for a given table entry are the same for most entries, especially when the SNR is large and the intrinsic EA is small. This is an indication that their respective pdfs are symmetric, and their values of $s$ and $\chi_{o}$ are consistent with the condition $s>3/\cos(2\chi_{o})$. The magnitude of the errors for the individual entries in the second column and last two rows of the table generally are not equal, indicating their respective pdfs are not symmetric.
Table 1 can serve as a lookup table for placing upper and lower limits on an EA measurement. Assuming that the measured EA is the mean EA, the table entries and the measured values of $s$ provide the errors associated with the measurement. For example, if the measured SNR is $s=3$ and the measured EA is $\chi_{m}=26.7^{\circ}$, the upper error on the measurement is $7.9^{\circ}$, and the lower error is $9.1^{\circ}$. Alternatively, analytical expressions for the upper and lower bounds as functions of $s$ for a given $\chi_{o}$ can be derived by fitting the data in Table 1 to equations similar to Equations A7 and A8. Or a more extensive lookup table with finer resolution in $s$ and $\chi_{o}$ can be generated, by numerically calculating the mean of the relevant pdf and its upper and lower bounds from Equation A6.